OpenStudy (anonymous):

x - y = 2 0.2x + 0.4y = 1 The x-coordinate of the solution to the system shown is _____. 1 3 6 got 1

OpenStudy (ikram002p):

got 3\[x-4=2\rightarrow4x+4y=10\] \[0.2x+0.4=1\rightarrow2x+4x=10\] \[6x=18\] \[x=3\]

OpenStudy (anonymous):

lol i kne wi did it wrong

OpenStudy (ikram002p):

mmm i wrot it fast :( \[x-y=2\rightarrow4x-4y=8\] now its better

OpenStudy (ikram002p):

\[0.2x+0.4y=1\rightarrow2x+4y=10\]

OpenStudy (anonymous):

kkk thanks

OpenStudy (ikram002p):

\[6x=18\]

OpenStudy (ikram002p):

you are wlc :)